OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1).

terenzreignz (terenzreignz):

To get the point-slope form, you need a slope, which is easily enough accomplished. Once you have the slope, choose any one of those two given points, and label it as (a,b). If m is the slope, then the point slope form of the equation is y - b = m(x - a)

terenzreignz (terenzreignz):

To get the slope, the formula is (y2 - y1)/(x2 - x1) But be careful; You must be consistent. If you decide to label, for instance, your y2, then its partner will be x2, and the same goes for x1 and y1 In this case, we let (x1, y1) = (-3, 2) (x2, y2) = (2, 1) [It can be the other way around, ie, choose (2,1) for your (x1, y1)] m = (1-2)/[2-(-3)] = -(1/5)

OpenStudy (anonymous):

I understand that, but whatever your giving me is not the right answer.

terenzreignz (terenzreignz):

My answer is y - 2 = -(1/5)(x + 3)

OpenStudy (anonymous):

Can you help me with this one Given the equation y − 4 = (x + 8) in point-slope form, identify the equation of the same line in slope-intercept form.

terenzreignz (terenzreignz):

slope intercept form is y = mx + b, try rearranging it to something like that by yourself first, and if you get stuck, just say so :)

OpenStudy (anonymous):

Oops , I typed it in wrong. It's y-4=3/4 (x+8) but I'll try first .

OpenStudy (anonymous):

Is it y=3/4x+12

terenzreignz (terenzreignz):

Where did you get 12?, did you forget to multiply 3/4 by 8?

OpenStudy (anonymous):

I don't get it

terenzreignz (terenzreignz):

y - 4 = (3/4)(x + 8) y - 4 = (3/4)x + (3/4)(8) y - 4 = (3/4)x + 6 y = (3/4)x + 6 + 4 y = (3/4)x + 10 //